# REPS: Recycled Entropy Packet Spraying for Adaptive Load Balancing and Failure Mitigation

### v3

@@ Abstract.
Next-generation datacenters require highly efficient network load balancing to manage the growing scale of artificial intelligence (AI) training and general datacenter traffic. Existing solutions designed for Ethernet, such as Equal Cost Multi-Path (ECMP) and oblivious packet spraying (OPS), struggle to maintain high network utilizations as datacenter topologies (and network failures as a consequence) continue to grow. To address these limitations, we propose REPS, a lightweight decentralized per-packet adaptive load balancing algorithm designed to optimize network utilization while ensuring rapid recovery from link failures. REPS adapts to network conditions by caching good-performing paths. In case of a network failure, REPS re-routes traffic away from it in less than 100 microseconds. REPS is designed to be deployed with next-generation out-of-order transports, such as Ultra Ethernet, and introduces less than 25 bytes of per-connection state. We extensively evaluate REPS in large-scale simulations and FPGA-based NICs.
copyright: nonedoi: isbn: price:
@@ 1. Introduction
Network architecture of distributed AI clusters is inherited from cloud workload deployments: WebSearch, Hadoop, Storage (Qian et al., 2024; Gangidi et al., 2024). It typically relies on Infiniband (269, 2004) and, more recently, commodity in-order Ethernet (e.g., RoCEv2 (Association, 2024)) for cost efficiency and ease of deployment. These solutions will struggle to deliver peak network bandwidth when scaled from $\approx 10$K training endpoints (e.g., to train models of GPT-4, Llama 3 size  (Dubey et al., 2024; Achiam et al., 2023)) to more than $100$K (e.g., next-generation model scales) due to:
@@
    management complexity and operational cost of lossless in-order network management at this scale due to link failures and degradation (Gangidi et al., 2024).
@@
Thus, the community recognized the need for network stacks tailored to distributed training traffic while remaining compatible with commodity datacenter Ethernet infrastructure (Gangidi et al., 2024). Such proposals include SDR by Amazon (Shalev et al., 2020), Falcon by Google (goo, 2023), TTPoE by Tesla (Motors, 2024), and the upcoming Ultra Ethernet (UE) (Consortium, 2024), developed in collaboration between major tech players. Key open questions in these proposals are how to address load balancing and mitigate link failures.
@@
Current-generation in-order Ethernet-based training systems (such as RoCEv2-based clusters) typically rely on ECMP (Hopps, 2009) or similar mechanisms for decentralized routing and load balancing. ECMP load balancing logic applies a hashing function to the 5-tuple header of each data packet to determine the next hop to take. The benefit of this scheme is that, ignoring link failures, it is unlikely to receive out-of-order packets at the destination NIC as the packets that belong to the same connection will be routed through the same network path.
@@
However, ECMP routing is fragile when different connections get hashed to the same link (Al-Fares et al., 2010; Alizadeh et al., 2014; Gangidi et al., 2024). In this scenario, flows can get hashed to the same path even when other paths are free, resulting in congestion and queue build up, which, in turn, can result in drops and go-back-N retransmission cycles (Hoefler et al., 2023).
@@
Moreover, recent works have shown that link failures drastically impact both training times and economic costs. A single link failure can have 20$\approx$ 20$\times$ 20 × higher cost impact in distributed training workloads than in cloud workloads(Qian et al., 2024; Gangidi et al., 2024). This observation, along with the increasing scale these systems are growing at, highlights the need for a transport layer with a load balancing scheme that can adapt near-instantly, e.g., within a few round-trip times (RTTs), to network link failures and, consequently, bandwidth asymmetries across the topology.
@@
Several solutions have been proposed to overcome ECMP’s limitations. MPTCP, PLB, FlowBender, Flowlet Switching, and Flowcell divide flows into subflows or flowlets and then route each one individually (Vanini et al., 2017; He et al., 2015; Kabbani et al., 2014, 2022; Raiciu et al., 2011). However, these solutions are still designed for in-order networks, making them inherently sensitive to collisions, prone to handling failures poorly, and requiring significant memory for each connection (Lu et al., 2018). Load balancers, such as Oblivious Packet Spraying (OPS) and Multi-Path RDMA (MPRDMA), which operate at a per-packet granularity, can mitigate ECMP-based collisions (Dixit et al., 2013; Lu et al., 2018). However, both approaches lack effective mechanisms to load balance effectively in the presence of failed network paths. Additionally, MPRDMA is constrained by its limited support for receiving out-of-order (OOO) packets and its requirement for per-packet acknowledgments (ACKs).
@@
Our key insight is that OPS problems can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as seen in SDR, UE, and Falcon (Lu et al., 2018; Dixit et al., 2013; NVIDIA, 2024; Consortium, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing scheme, Recycled Entropy Packet Spraying aka REPS. REPS caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
@@
REPS does not require any specific hardware support from switches beyond ECMP-like header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). REPS requires only $\approx 25$ bytes of state per-connection, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
@@
We extensively evaluate REPS in simulation and augment the simulation findings with real hardware results in Section 4. We deploy REPS in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, REPS consistently outperforms state-of-the-art load balancing algorithms. REPS outperforms ECMP and OPS by up to 6×6$\times$6 × and 1.25×1.25$\times$1.25 × in symmetric networks, and by up to 4.5×4.5$\times$4.5 × and 1.5×1.5$\times$1.5 × in asymmetric networks, outperforming OPS by as much as 100×100$\times$100 × during short-term transient link failures.
2. Background
@@
In this section we introduce the different building blocks necessary to understand REPS logic. We first describe different congestion signals (ECN, loss). We then introduce key terms and concepts related to load balancing (ECMP, EV).
@@ 2.1. Congestion Signals
allows switches to notify congestion by setting a bit in the traffic class field of the IP header. The receiver then sends back this marked ECN bit to the sender in its ACK packet header. The sender can choose to react to this congestion signal by adjusting its sending rate (Zhu et al., 2015; Alizadeh et al., 2010a; Ye et al., 2019). Switches can employ various strategies for marking packets. For instance, in Random Early Detection (RED)(Floyd and Jacobson, 1993), switches probabilistically mark packets based on queue size, with the marking probability increasing linearly between two thresholds ($K_{min}$ and $K_{max}$).
In REPS, we use ECN as the congestion signal to detect path congestion due to its simplicity and widespread adoption (Hoefler et al., 2023; Nvidia, 2024). Since ECN is not marked for packets when the queue is smaller than $K_{min}$, ECN effectively filters out cases of minor queuing due to packet collisions across multiple hops, while identifying true congestion at a single bottleneck. In contrast, delay-based signals struggle to differentiate between these scenarios unless enhanced by advanced switch features, such as in-network telemetry (INT) (Wang et al., 2023).
@@
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two type of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). REPS can optionally use trimming, if supported, to distinguish congestion losses from network failure, which are indicated by timeouts (Appendix A).
@@
rely on congestion signals, such as ECN marks or packet loss, to adjust a flow’s sending rate or window with the goal of maintaining high network utilization while preventing queue build up. REPS is designed to work well with any CC algorithm as long as they support receiving and acknowledging packets out-of-order for a given message. In particular we show later how REPS works well with EQDS, a variant of DCTCP, and an internal CC algorithm (Alizadeh et al., 2010b; Olteanu et al., 2022).
@@ 2.2. Load Balancing
Under normal circumstances, all packets belonging to the same flow are assigned statically to a given path since the hashing function will use the same values as input. This assignment is done statically and ignores the current network congestion and failure conditions. As a result, two or more flows might be assigned to the same path even if there are many more paths available. This will inevitably result in heavy congestion and possibly packet drops as a consequence. Such ECMP hash collisions are a well-documented limitation of standard ECMP (Gangidi et al., 2024; Alizadeh et al., 2014; Zhang et al., 2021).
@@
is a value in the packet header that can be configured to be an input to the hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). We leverage EVs in REPS to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
@@
is a fixed-size set of EVs, since the number of possible values is constrained by the number of bits they can occupy in the packet header. For example, the source port field in a UDP header is assigned 16 bits, giving the EVS a size of 65536 possible values (excluding some reserved values) (Lu et al., 2018). While different numbers of bits could be allocated for the EVS, we analyze in Section 4.5.1 and Appendix B how many are required for optimal performance. It is generally advantageous for an algorithm to achieve good load balancing performance with a small EVS size since that often reduces the algorithm’s memory overhead.
@@
3. REPS
Recycled Entropy Packet Spraying aka REPS is a load balancing algorithm that relies on simple and memory-efficient endpoint logic. By design, REPS can be implemented in NIC hardware or firmware with minimal memory/area footprint. REPS does not need any switch support besides ECMP hashing and ECN marking. The key idea behind REPS is straightforward: when congestion is detected on a certain path, we explore alternative paths while caching and reusing paths with little to no congestion. Specifically, REPS uses a circular buffer of a fixed size to cache EVs of uncongested paths. Algorithm 1 details the pseudocode for a REPS sender when receiving an ACK and when detecting a failure, and Algorithm 2 describes the pseudocode for a REPS sender when sending out a data packet.
@@ 3.1. Core Logic: Path Exploration and Reuse
During the first Bandwidth-Delay Product (BDP) packets of a new flow, a REPS sender explores random entropies from the EVS. This exploration is necessary because, initially, there is no knowledge about the network’s state. In this warm up phase, REPS operates similarly to OPS.
@@
When an ACK arrives at the sender, if it is not ECN-marked, the EV it carries is cached in the circular buffer, and its validity bit is set to 1. Otherwise, if the ACK is ECN-marked, REPS does not cache the EV and discards it. When set, the validity bit indicates that an entropy has not been used after it has been added to the the circular buffer. When sending a data packet out, REPS first checks if there are any valid EVs in its buffer. In case there is any valid EV, REPS reuses the oldest valid EV from the circular buffer and resets its validity bit. Otherwise, REPS explores a random EV from the EVS.
The circular buffer in REPS ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
Once the network experiences any kind of transient (e.g. link flap) or persistent unrecoverable failure (e.g., a link or switch failure), it will take the system some time to recover from it: ranging from several milliseconds to update the ECMP routing group to several seconds if a reboot is needed, and much more if a swap is needed (Anbiah and Sivalingam, 2022; Iselt et al., 2004).
@@
REPS detects such failures via indirect feedback from the network and enters freezing mode. REPS uses a simple timeout heuristic (Section 2.1) that can be enhanced with packet trimming (Appendix A) as a natural feedback from the network to detect failures along a path.
When in freezing mode, REPS:
@@
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that REPS will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we set a timer which can be configured by the operator. For instance, in our internal testing with real hardware, we have found that this can be equal to the maximum observed time that it takes for a failure to recover plus a certain buffering period. Once we exit freezing mode, we use random EVs to allow REPS to explore new paths and assess whether we detect new packets failing or not. This prevents REPS from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a dead path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if REPS exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (Figure 7 and Figure 8).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), REPS would still load balance well, as discussed in Section 4.5.1 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, REPS can be conservative and can safely enter freezing mode.
Algorithm 1 REPS logic upon ACK receive and failure detection.
@@
5:procedure onAck(ackPacket)
@@
9:    if not repsBuffer[head].isValid then
@@
15:    if isFreezingMode and now() ¿ exitFreezingMode then
@@
Algorithm 2 REPS logic on send datapath.
@@
15:    if repsBuffer.isEmpty() or (numberValidEVs == 0 and not isFreezingMode) or exploreCounter then
@@
19:        $dataPacket.ev=\textit{getNextEntropy()}$()
@@
3.3. REPS Design Advantages

### v4

# ARCANE: Adaptive Routing with Caching and Aware Network Exploration

@@ Abstract.
Next-generation datacenters require highly efficient network load balancing to manage the growing scale of artificial intelligence (AI) training and general datacenter traffic. However, existing Ethernet-based solutions, such as Equal Cost Multi-Path (ECMP) and oblivious packet spraying (OPS), struggle to maintain high network utilization due to both increasing traffic demands and the expanding scale of datacenter topologies, which also exacerbate network failures. To address these limitations, we propose ARCANE, a lightweight decentralized per-packet adaptive load balancing algorithm designed to optimize network utilization while ensuring rapid recovery from link failures. ARCANE adapts to network conditions by caching good-performing paths. In case of a network failure, ARCANE re-routes traffic away from it in less than 100 microseconds. ARCANE is designed to be deployed with next-generation out-of-order transports, such as Ultra Ethernet, and uses less than 25 bytes of per-connection state regardless of the topology size. We extensively evaluate ARCANE in large-scale simulations and FPGA-based NICs.
submissionid: 52copyright: nonedoi: isbn: price:
@@ 1. Introduction
Modern AI training cluster networks draw on both high-performance computing (HPC) and cloud-native architectures (Hoefler et al., 2022b; Qian et al., 2024; Gangidi et al., 2024). These systems rely on Remote Direct Memory Access (RDMA) for low-latency, high-throughput communication. InfiniBand, a specialized RDMA transport, is widely used for its performance (269, 2004). More recently, operators have adopted commodity Ethernet fabrics, especially RoCEv2 (Association, 2024), which enable RDMA over standard hardware using off-the-shelf switches and NICs to reduce costs. However, as training clusters scale from around 10,000 endpoints (enough to train GPT-4 or Llama 3–scale models (Achiam et al., 2023; Dubey et al., 2024)) to over 100,000 nodes, both interconnects face fundamental limits in sustaining peak bandwidth due to:
@@
    the complexity and overhead of managing a lossless, in-order network at scale, especially under link failures and performance degradation. (Gangidi et al., 2024).
@@
Thus, the community recognized the need for network stacks tailored to distributed training traffic while remaining compatible with commodity datacenter Ethernet infrastructure (Gangidi et al., 2024). Such proposals include SRD by Amazon (Shalev et al., 2020), Falcon by Google (goo, 2023), TTPoE by Tesla (Motors, 2024), and the upcoming Ultra Ethernet (UE) (Consortium, 2024), developed in collaboration between major tech players. Key open questions in these proposals are how to address load balancing and mitigate link failures.
@@
Current-generation in-order Ethernet-based training systems (such as RoCEv2-based clusters) typically rely on ECMP (Hopps, 2009) or similar mechanisms for decentralized routing and load balancing. ECMP load balancing logic applies a hashing function to the 5-tuple header of each data packet to determine the next hop to take. The benefit of this scheme is that, ignoring link failures, it is unlikely to receive out-of-order packets at the destination NIC as packets belonging to the same connection will be routed through the same network path.
@@
However, ECMP routing is fragile when different connections get hashed to the same link (Al-Fares et al., 2010; Alizadeh et al., 2014; Gangidi et al., 2024). In this scenario, flows can get hashed to the same path even when other paths are free, resulting in congestion and queue build up, which, in turn, can result in drops and go-back-N retransmissions (Hoefler et al., 2023).
@@
Moreover, recent works have shown that link failures drastically impact both training times and economic costs. A single link failure can have 20$\approx$ 20$\times$ 20 × higher cost impact in distributed training workloads than in cloud workloads (Qian et al., 2024; Gangidi et al., 2024). This observation, along with the increasing scale these systems are growing at, highlights the need for a transport layer with a load balancing scheme that can adapt near-instantly, e.g., within a few round-trip times (RTTs), to network failures and, consequently, bandwidth asymmetries across the topology.
@@
Several solutions have been proposed to overcome ECMP’s limitations. MPTCP (Raiciu et al., 2011), PLB (Qureshi et al., 2022), FlowBender (Kabbani et al., 2014), Flowlet Switching (Vanini et al., 2017), and Flowcell (He et al., 2015) divide flows into subflows or flowlets and then route each one individually. However, these solutions are still designed for in-order networks, making them inherently sensitive to collisions, prone to handling failures poorly, and requiring significant memory for each connection (Lu et al., 2018). Load balancers, such as Oblivious Packet Spraying (OPS) (Dixit et al., 2013) and Multi-Path RDMA (MPRDMA)  (Lu et al., 2018), that operate at a per-packet granularity, can mitigate ECMP-based collisions. However, both approaches lack mechanisms to load balance effectively in the presence of network failures. Additionally, MPRDMA is constrained by its limited support for receiving out-of-order (OOO) packets and its requirement for per-packet acknowledgments (ACKs).
@@
Our key insight is that the challenges of spraying-based load balancing can be addressed by adaptive packet spraying, paired with a transport layer that natively supports out-of-order packet delivery, as demonstrated in SDR (Shalev et al., 2020), UE (Lu et al., 2018), Falcon (goo, 2023) and recent research work (Dixit et al., 2013; NVIDIA, 2024; Gangidi et al., 2024). Based on this insight, we design and contribute to the Ultra Ethernet Consortium (UEC) our decentralized load-balancing algorithm, Adaptive Routing with Caching and Aware Network Exploration aka ARCANE. ARCANE caches ”good” network paths in a circular buffer and quickly recovers (within a few RTTs) from network failures by adaptively discovering or freezing network paths.
@@
ARCANE does not require any specific hardware support from switches beyond ECMP-based header hashing and ECN, which are standard features in modern switches (Hoefler et al., 2023; Nvidia, 2024). ARCANE requires only $\approx 25$ bytes of state per-connection regardless of the topology size, whereas MPTCP requires 368 extra bytes for 8 sub-flows (Lu et al., 2018).
@@
We extensively evaluate ARCANE in simulation and augment the simulation findings with real hardware results in Section 4. We deploy ARCANE in a cluster with modified FPGA-based RDMA-capable NICs. In large-scale network simulations, ARCANE consistently outperforms state-of-the-art load balancing algorithms. ARCANE outperforms ECMP and OPS by up to 6×6$\times$6 × and 1.25×1.25$\times$1.25 × in symmetric networks, and by up to 4.5×4.5$\times$4.5 × and 2×2$\times$2 × in asymmetric networks, outperforming OPS by as much as 100×100$\times$100 × during short-term transient link failures.
2. ARCANE Building Blocks
@@
In this section, we introduce the fundamental building blocks necessary to understand ARCANE logic. We first describe key congestion signals, such as ECN and packet loss. We then introduce essential terms and concepts related to load balancing, including ECMP and EV. These concepts are largely established, but we revisit them here to provide clear terminology and a basis for the discussion that follows.
@@ 2.1. Congestion Signals
allows switches to notify congestion by setting a bit in the traffic class field of the IP header. The receiver then sends back this marked ECN bit to the sender in its ACK packet header. The sender can choose to react to this congestion signal by adjusting its sending rate (Zhu et al., 2015; Alizadeh et al., 2010; Ye et al., 2019). Switches can employ various strategies for marking packets. For instance, in Random Early Detection (RED) (Floyd and Jacobson, 1993), switches probabilistically mark packets based on queue size, with the marking probability increasing linearly between two thresholds ($K_{min}$ and $K_{max}$).
In ARCANE, we use ECN as the congestion signal to detect path congestion due to its simplicity and widespread adoption (Hoefler et al., 2023; Nvidia, 2024). Since ECN is not marked for packets when the queue is smaller than $K_{min}$, ECN effectively filters out cases of minor queuing due to packet collisions across multiple hops, while identifying true congestion at a single bottleneck. In contrast, delay-based signals struggle to differentiate between these scenarios unless enhanced by advanced switch features, such as in-network telemetry (INT) (Wang et al., 2023).
@@
has long been a key indicator of severe congestion in networks (Floyd and Jacobson, 1993; Nichols and Jacobson, 2012). However, using packet losses as the sole signal for congestion detection can result in delayed responses, as losses typically indicate a point of significant congestion. Packet loss detection, often based on timeouts, can be challenging to calibrate and may lead to unnecessary retransmissions. We categorize packet losses in two categories: losses because of severe congestion and losses due to networking failures. To differentiate the two types of losses and to improve reaction times, packet trimming, which only triggers for congestion drops, can be employed  (Adrian et al., 2022; Cheng et al., 2014; Handley et al., 2017; Olteanu et al., 2022). Trimming can be implemented on many existing switches (Adrian et al., 2022) and is starting to be supported by some switch vendors. It is also being pushed as a UE protocol feature (Consortium, [n. d.]). On current generation networks, ARCANE can be deployed with timeouts, and in next-generation augmented with trimming to distinguish between loss types (more details in Appendix A).
@@
rely on congestion signals, such as ECN marks or packet loss, to adjust a flow’s sending rate or window with the goal of maintaining high network utilization while preventing queue build up. ARCANE is designed to work well with any CC algorithm as long as they support receiving and acknowledging packets out-of-order for a given message. In particular, we show later how ARCANE works well with EQDS, a variant of DCTCP, and a proprietary internal CC algorithm (Olteanu et al., 2022; Alizadeh et al., 2010).
@@ 2.2. Load Balancing
Under normal circumstances, all packets belonging to the same flow are assigned statically to a given path since the hashing function will use the same values as input. This assignment is done statically and ignores the current network congestion and failure conditions. As a result, two or more flows might be assigned to the same path even if there are many more paths available. This will inevitably result in heavy congestion and, possibly, packet drops as a consequence. Such ECMP hash collisions are a well-documented limitation of standard ECMP (Gangidi et al., 2024; Alizadeh et al., 2014; Zhang et al., 2021).
@@
is a value in the packet header that can be configured to be an input to the ECMP hashing function in switches. Such a value, which is set by the sender, allows it to alter a packet’s path in the network. Possible header fields to be used as an EV are the Source Port field in the packet header (Lu et al., 2018) or the Flow Label field in IPv6 (Qureshi et al., 2022). This is consistent with previous load balancing works and has already been demonstrated in practice (Kabbani et al., 2014; Qureshi et al., 2022; Vanini et al., 2017; Lu et al., 2018). We leverage EVs in ARCANE to improve load balancing and address the ECMP collision limitation, without needing to know the exact mapping between a packet’s EV and its resulting path.
@@
is a fixed-size set of EVs, since the number of possible values is constrained by the number of bits they can occupy in the packet header. For example, the source port field in a UDP header is assigned 16 bits, giving the EVS a size of 65536 possible values (excluding some reserved values) (Lu et al., 2018). While different numbers of bits could be allocated for the EVS, we analyze in Section 4.5.2 and Appendix B how many are required for optimal performance. It is generally advantageous for an algorithm to achieve good load balancing performance with a small EVS size since that often reduces the algorithm’s memory overhead. Note that the minimum effective EVS size also depends on the network topology: smaller topologies require fewer EVs to fully explore all available physical paths.
@@
3. ARCANE
ARCANE is a load balancing algorithm that relies on simple and memory-efficient endpoint logic. The key idea behind ARCANE is straightforward: when congestion is detected on a certain path, we explore alternative paths while caching and reusing paths with little to no congestion. Specifically, ARCANE uses a circular buffer of a fixed size to cache EVs of uncongested paths. By design, ARCANE can be implemented in NIC hardware or firmware with minimal memory/area footprint. ARCANE does not need any switch support besides ECMP hashing and ECN marking. Placing ARCANE at the endpoint not only simplifies deployment but also provides a broader, end-to-end view of the network, unlike fabric-side approaches which are limited to local visibility. Algorithm 1 details pseudocode for a ARCANE sender receiving an ACK and detecting a failure. Algorithm 2 describes pseudocode for a ARCANE sender sending a data packet.
@@ 3.1. Core Logic: Path Exploration and Reuse
During the first Bandwidth-Delay Product (BDP) worth of packets of a new or idle connection, a ARCANE sender explores random entropies from the EVS. This exploration is necessary because, initially, there is no fresh knowledge about the network’s state. In this warm up phase, ARCANE operates similarly to OPS.
@@
When an ACK arrives at the sender, if it is not ECN-marked, the EV it carries is cached in the circular buffer, and its validity bit is set to 1. Otherwise, if the ACK is ECN-marked, ARCANE does not cache the EV and discards it. When set, the validity bit indicates that an entropy has not been used after it has been added to the circular buffer. When sending a data packet out, ARCANE first checks if there are any valid EVs in its buffer. In case there is any valid EV, ARCANE reuses the oldest valid EV from the circular buffer and resets its validity bit. Otherwise, ARCANE explores a random EV from the EVS.
The circular buffer in ARCANE ensures that bursts of back-to-back ACKs with ”good” entropies are correctly cached and reused. Moreover, it guarantees stable load balancing in the case of failures as shown in Section 3.2. We use a circular buffer of 8 elements based on the bounds from Theorem 5.1.
@@ 3.2. Failure Mitigation: Freezing Mode
When a network, running without ARCANE, experiences any kind of transient (e.g., link flap) or persistent unrecoverable failure (e.g., a link or switch failure), it will take the system some time to recover from it: ranging from several milliseconds to update the ECMP routing group to several seconds if a reboot is needed, and more if a swap is needed (Anbiah and Sivalingam, 2022; Iselt et al., 2004).
@@
By default, ARCANE Core Logic would already limit utilizing failing paths since it would only recycle paths that are still active and returning ACKs without ECN. However, if the ARCANE buffer was empty, it would select a random EV that could potentially map to a failing path. To further improve the performance of ARCANE and prevent further losses, we introduce freezing mode. ARCANE enters this mode when detecting failures via indirect feedback from the network. ARCANE uses a timeout heuristic (Section 2.1) that can be enhanced with packet trimming (Appendix A) as a natural feedback from the network to detect failures along a path.
When in freezing mode, ARCANE:
@@
While this strategy could result in slightly worse load balancing (due to potentially reusing the same EV several times), it comes with the major benefit of guaranteeing that ARCANE will almost never pick the failing path again since the recent received EVs point to healthy paths. Considering the example above: by enabling freezing mode, the number of packets dropped decreases from over 120K packets to only about 1K.
To decide when to exit freezing mode, we either send probe packets to check the status of the failing paths or, if probing is not available, we simply exit freezing mode after a fixed amount of time. Once we exit freezing mode, we use random EVs to allow ARCANE to explore new paths and assess whether we detect new packets failing or not. This prevents ARCANE from getting stuck in a suboptimal state if the EVs in the buffer were all pointing to a failing path, a rare scenario that can theoretically happen with properly timed back-to-back network failures. Moreover, if ARCANE exits freezing mode before the issue is fully resolved, it will simply re-enter the mode shortly afterward with minimal impact on performance (as we will show in Section 4.3.3).
The intuition behind freezing mode is that once we suspect there is a failure, we want to start avoiding it as soon as possible. Interestingly, we observe that even if we enter freezing mode unnecessarily (i.e., by mistaking a congestion drop for a network failure), ARCANE would still load balance well, as discussed in Section 4.5.2 and Appendix A. This observation means that even if there is doubt about whether a real failure occurred, ARCANE can be conservative and can safely enter freezing mode.
Algorithm 1 ARCANE logic upon ACK receive and failure detection.
@@
5:procedure onAck(ackPacket) $\triangleright$ Sec. 3.1
@@
9:    if not ARCANEBuffer[head].isValid then
@@
15:    if isFreezingMode and now() ¿ exitFreezingMode then $\triangleright$ Sec. 3.2
@@
Algorithm 2 ARCANE logic on send datapath.
@@
15:    if ARCANEBuffer.isEmpty() or (numberOfValidEVs == 0 and not isFreezingMode) or exploreCounter then
@@
19:        $dataPacket.ev=\textit{getNextEV()}$()
@@
3.3. ARCANE Design Advantages